A truck runs into a pile of sand, moving 1.6 m as it slows to a stop. The magnitude of the work that the sand does on the truck is 6.0×105J.

Determine the magnitude of the average force that the sand exerts on the truck. Did the sand do positive or negative work? How does the magnitude of the average force change if the stopping distance is doubled? Assume that the work that the sand does on the truck remains unchanged.
Answer #1

Work = force * perpendicular distance

W = F*d

6*10^5 = F*1.6

F = 3.75*10^5 N

if the stopping distance is doubled , The force will be halved